OpenStudy (anonymous):

newtons second law of motion if the force on an object is 20N and its mass 5 g, what is its acceleration?

OpenStudy (anonymous):

F=ma, right? So you have F and m, what is a?

OpenStudy (anonymous):

20=5g

OpenStudy (anonymous):

doesn't give it, that is the whole question

OpenStudy (anonymous):

F=ma, F=20N, m=5g, so you have \(20N=5g\times a\), solve for \(a\).

OpenStudy (anonymous):

I have looked on 2nd law but so confusing

OpenStudy (anonymous):

4

OpenStudy (anonymous):

4??

OpenStudy (anonymous):

There is actually one other trick here that I hadn't noticed. Newtons are kilogram meters per second squared, but it's giving you mass in grams, not kilograms. So you have to adjust that.

OpenStudy (anonymous):

.005kg?

OpenStudy (anonymous):

So, 4g is .004kg. So you have 20kgm/s^2=.005kg*a m/s^2. So a=20/.005m/s^2, which is 4,000 m/s^2.

OpenStudy (anonymous):

Sorry, 5g is .005kg that should read.

OpenStudy (anonymous):

ty so much